# Ensembling Portfolio Strategies for Long-Term Investments: A Distribution-Free Preference Framework for Decision-Making and Algorithms

### v1

@@ 4.3 Experiments with large scales of component strategies
After conducting eight experiments in the previous section, we further extend to the UC strategy, given by (3.3), and its accelerated variant in a scenario involving a large scale of component M-V strategies. For this purpose of our analysis, we consider a set $\mathcal{A}$ of risk aversion coefficients that encompasses values ranging from $0.001$ to $0.009$ with a step size of $0.001$, from $0.01$ to $0.09$ with a step size of $0.01$, from $0.1$ to $0.9$ with a step size of $0.1$, and from $1$ to $10$ with a step size of $0.5$. In Experiment 9, we examine the effects of partitioning $\mathcal{A}$ into multiple base sets, along with their respective sizes, on the performance and behavior of the UC strategy. To investigate this further, we analyze the impact of partitions into two, four, and six base sets. More specifically, in the context of two base sets, our analysis includes examining balanced sets with equal sizes spanning from $0.001$ to $0.5$ and from $0.6$ to $10$, respectively; as well as considering two unbalanced base sets covering the ranges of $0.001$ to $0.02$ and $0.03$ to $10$, respectively. On the other hand, for the case of four base sets, we divide the coefficients from $0.001$ to $5$ into three equal-sized base sets, while assigning the remainder to the fourth base set. In a similar vein, when dealing with six base sets, we evenly segment the coefficients ranging from $0.001$ to $7$ into the first five base sets, while leaving the remainder for the final one. Following Experiment 9, we proceed to implement accelerated strategies tailored to the scenario involving six base sets in Experiment 10.
Refer to captionRefer to caption
@@
Performance measures of the UC strategy, along with a WAE strategy applied to all component M-V strategies, are presented in Figure 10 and Table 3. The results demonstrate that the WAE strategy performs poorly in comparison to the UC strategy across all partitions, despite its higher Sharpe ratio. This is due to the growth rate of the WAE strategy being bounded by that of the best component M-V strategy. Overall, all partitions lead to the outperformance of the UC strategy over the best M-V strategy in terms of cumulative wealth. However, the partition into unbalanced base sets shows relatively lower cumulative wealth compared to more evenly segmented base sets. Furthermore, increasing the number of base sets while reducing the size of each set simultaneously improves both the average growth rate, implying higher final cumulative wealth, and the average return, but results in a lower Sharpe ratio. These findings align with the characteristics observed in the previous eight experiments of the UC strategy. In more detail, regarding cumulative wealth, the UC strategy with two unbalanced base sets merely matches the cumulative wealth of the best M-V strategy after approximately $6600$ time periods, while the UC strategy with two balanced base sets eventually exceeds it. After the same amount of time, the UC strategies with four and six base sets exceed the best M-V strategy, with slightly more favorable cumulative wealth for the latter partitioning. This distinction is clearly depicted in Figure 11.
In the first row of Figure 11, it is evident that a higher number of base sets leads to a slower convergence speed for the growth rates of the benchmark and the UC strategies. Conversely, partitioning into smaller-sized base sets can rapidly lift the difference $W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}$ above zero over time, enabling the UC strategy to exceed the baseline strategy faster. However, this difference may be impeded from exceeding zero by partitioning into strongly unbalanced base sets, as observed in the first upper graph. This phenomenon is caused by the behaviors of the representative strategies of the respective base sets, as a base set may include numerous M-V strategies with markedly different behaviors from the others in the same set. Additionally, the second row of Figure 11 demonstrates the upper bound for the difference $W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{\text{UC}}\big{)}$ established in Proposition 5, with respect to the numerically approximated UC strategy. It shows that, in general, the average difference over time converges to zero slower when the number of base sets is higher. The results of Experiment 9 highlight the importance of appropriately tuning the partition for the set $\mathcal{A}$, such that the number of base sets should not be large, while each base set should mainly contain component strategies that behave in a similar manner.
@@
In Experiment 10, we conduct the UC-L and UC-W strategies alongside the UC strategy using a partition of six base sets. The computation procedure for the UC-L and UC-W strategies remains unchanged, except for the replacement of component M-V strategies with representative strategies. The results of this experiment are presented in Figure 11 and Table 3. Similar to the case of a small scale of the component strategies, the UC-L strategy with smaller sets $\bar{B}_{n}$ exhibits better cumulative wealth and average return compared to the other two, albeit with a slight decrease in the Sharpe ratio. Figure 12 illustrates the allocation proportions of capital to each component M-V strategy over time by the UC-L ($30\%$) strategy, in comparison to the original UC strategy with all partitions except the unbalanced one. The graphs display that the proportions experience lower variability as the number of base sets increases, while demonstrating stronger fluctuations in the presence of acceleration for the UC strategy. This phenomenon can be explained using the results from the experiments conducted with small scales of risk aversion levels. For a smaller number of base sets, a larger number of component strategies are adjusted relative to the weights assigned to the corresponding representative strategies, and vice versa. Moreover, the changes in weights on each representative strategy are more gradual when their number is higher, as depicted in the second row of Figure 7. Therefore, the extent of fluctuation in the proportions of the UC strategy decreases as the number of base sets increases. However, since acceleration causes strong variations in the weights of representative strategies, the proportions of the component strategies also exhibit wild fluctuations, as depicted in the last graph for the UC-L strategy.
@@ 5 Summary and concluding Remarks
$\forall\epsilon>0,\,\exists N_{\epsilon}\geq 1\text{ such that }\sup_{n\geq N}\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}\big{)}\big{)}<\epsilon,\,\forall N\geq N_{\epsilon}.$
@@
where the vectors $\big{(}\big{<}b_{t}^{1},x_{t}\big{>},...,\big{<}b_{t}^{k},x_{t}\big{>}\big{)}\eqqcolon\big{(}r_{t,1},...,r_{t,k}\big{)}\eqqcolon r_{t}\in\mathbb{R}_{+}^{k}$ represent the respective returns of $k$ component portfolios at time $t$. Let $r_{1}^{n}\coloneqq\big{\{}r_{1},...,r_{n}\big{\}}\in\mathbb{R}_{+}^{k\times n}$ denote the sequence of vectors of portfolios’ returns. Then, by applying the theorem of the upper bound for the worst-case of the Universal portfolio in Cover (1991); Cover and Ordentlich (1996), with replacing the sequence of assets’ returns by $r_{1}^{n}$, we obtain the upper bound with uniform density $\mu\left(\lambda\right)$ as:
@@
Thus, ${\displaystyle\max_{x_{1}^{n}}\left(\log S_{n}\left(b_{n}^{*}\right)-\log S_{n}\left(b_{n}\right)\right)\leq\left(k-1\right)\log\left(n+1\right)}$ for all $n$, as needed to show
@@
Then, we have the following upper bound for every representative strategy:
@@
On the other hand, by Proposition 4 and the property of the benchmark strategy for this case as $S_{n}\big{(}b_{n}^{*}\big{)}\geq\max_{i\in\left[N\right]}S_{n}\big{(}\hat{b}_{n}^{i}\big{)}$ for all $n$, we obtain the needed upper bound as follows:
@@
The last inequality is due to (A.1) and the fact that $S_{n}\big{(}\ddot{b}_{n}\big{)}=\max_{i\in\left[N\right]}S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}$.

### v2

@@ 4.3 Experiments with large scales of component strategies
After conducting eight experiments in the previous section, we extend our investigation to the UC strategy given by (3.3) and its accelerated variant in a scenario involving a large scale of component M-V strategies. For this analysis, we consider a set $\mathcal{A}$ of risk aversion coefficients encompassing values ranging from $0.001$ to $0.009$ with a step size of $0.001$, from $0.01$ to $0.09$ with a step size of $0.01$, from $0.1$ to $0.9$ with a step size of $0.1$, and from $1$ to $10$ with a step size of $0.5$. In Experiment 9, we examine the impact of partitioning $\mathcal{A}$ into multiple base sets and their respective sizes on the performance and behavior of the UC strategy. To investigate this further, we analyze partitions into two, four, and six base sets. Specifically, in the case of two base sets, our analysis includes examining balanced sets with equal sizes spanning from $0.001$ to $0.5$ and from $0.6$ to $10$, as well as unbalanced sets covering the ranges of $0.001$ to $0.02$ and $0.03$ to $10$. In the case of four base sets, we divide the coefficients from $0.001$ to $5$ into three equal-sized sets, while assigning the remainder to the fourth set. In a similar vein, for six base sets, we evenly segment the coefficients ranging from $0.001$ to $7$ into the first five sets, while leaving the remainder for the final set. Following Experiment 9, we proceed to implement accelerated strategies tailored to the scenario involving six base sets in Experiment 10.
@@
Performance measures of the UC strategy, along with a WAE strategy applied to all component M-V strategies, are presented in Figure 10 and Table 3. The results demonstrate that the WAE strategy performs poorly in comparison to the UC strategy across all partitions, despite its higher Sharpe ratio. This is because the growth rate of the WAE strategy is bounded by that of the best component M-V strategy. Overall, all partitions lead to the outperformance of the UC strategy over the best M-V strategy in terms of cumulative wealth. However, the partition into unbalanced base sets shows relatively lower cumulative wealth compared to more evenly segmented base sets. Furthermore, increasing the number of base sets while reducing the size of each set simultaneously improves both the average growth rate, implying higher final cumulative wealth, and the average return, but results in a lower Sharpe ratio. These findings align with the characteristics observed in the previous eight experiments of the UC strategy. In more detail, regarding cumulative wealth, the UC strategy with two unbalanced base sets merely matches the cumulative wealth of the best M-V strategy after approximately $6600$ time periods, while the UC strategy with two balanced base sets eventually exceeds it. After the same amount of time, the UC strategies with four and six base sets exceed the best M-V strategy, with slightly more favorable cumulative wealth for the latter partitioning. This distinction is clearly depicted in Figure 11.
In the first row of Figure 11, it is evident that a higher number of base sets leads to a slower convergence speed for the growth rates of the benchmark and UC strategies. Conversely, partitioning into smaller-sized base sets can rapidly lift the difference $W_{n}\big{(}b_{n}^{*}\big{)}-W_{n}\big{(}\ddot{b}_{n}\big{)}$ above zero over time, enabling the UC strategy to exceed the baseline strategy faster. However, this difference may be impeded from exceeding zero by partitioning into strongly unbalanced base sets, as observed in the first upper graph. This phenomenon is caused by the behaviors of the representative strategies of the respective base sets, as a base set may contain numerous M-V strategies with markedly different behaviors from the others in the same set. Additionally, the second row of Figure 11 illustrates the upper bound for the difference $W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}^{\text{UC}}\big{)}$, established in Proposition 5, with respect to the numerically approximated UC strategy. It shows that, in general, the average difference over time converges to zero more slowly when the number of base sets is higher. The results of Experiment 9 emphasize the importance of appropriately tuning the partition for the set $\mathcal{A}$, such that the number of base sets should not be large, while each base set should mainly contain component strategies that behave in a similar manner.
@@
In Experiment 10, we conduct the UC-L and UC-W strategies alongside the UC strategy using a partition of six base sets. The computation procedure for the UC-L and UC-W strategies remains unchanged, except for the replacement of component M-V strategies with representative strategies. The results of this experiment are presented in Figure 11 and Table 3. Similar to the case with a small scale of component strategies, the UC-L strategy with a small set $\bar{B}_{n}$ exhibits higher cumulative wealth and average return compared to the other two, albeit with a slight decrease in the Sharpe ratio. Figure 12 illustrates the allocation proportions of capital to each component M-V strategy over time by the UC-L ($30\%$) strategy, in comparison to the original UC strategy with all partitions except the unbalanced one. The graphs show that the proportions experience lower variability as the number of base sets increases, while demonstrating stronger fluctuations under acceleration for the UC strategy. This phenomenon can be explained using the results from the experiments conducted with small scales of risk aversion levels. Specifically, with fewer base sets, a larger number of component strategies are adjusted relative to the weights assigned to their corresponding representative strategies, and vice versa. Moreover, changes in the weights of each representative strategy are more gradual when their number is higher, as depicted in the second row of Figure 7. Therefore, the extent of fluctuation in the proportions of the UC strategy decreases as the number of base sets increases. However, since acceleration causes strong variations in the weights of representative strategies, the proportions of the component strategies also exhibit significant fluctuations, as shown in the last graph for the UC-L strategy.
@@ 5 Summary and concluding Remarks
$\forall\epsilon>0,\,\exists N_{\epsilon}\text{ such that }\sup_{n\geq N}\big{(}W_{n}\big{(}\ddot{b}_{n}\big{)}-W_{n}\big{(}b_{n}\big{)}\big{)}<\epsilon,\,\forall N\geq N_{\epsilon}.$
@@
where the vectors $\big{(}\big{<}b_{t}^{1},x_{t}\big{>},...,\big{<}b_{t}^{k},x_{t}\big{>}\big{)}\eqqcolon\big{(}r_{t,1},...,r_{t,k}\big{)}\eqqcolon r_{t}\in\mathbb{R}_{++}^{k}$ represent the respective returns of $k$ component portfolios at time $t$. Let $r_{1}^{n}\coloneqq\big{\{}r_{1},...,r_{n}\big{\}}\in\mathbb{R}_{++}^{k\times n}$ denote the sequence of $n$ vectors of portfolios’ returns. Then, by applying the theorem of the upper bound for the worst-case of the Universal portfolio in Cover (1991); Cover and Ordentlich (1996), with replacing the sequence of assets’ returns by $r_{1}^{n}$, we obtain the upper bound with uniform density $\mu\left(\lambda\right)$ as:
@@
Thus, ${\displaystyle\max_{x_{1}^{n}}\left(\log S_{n}\left(b_{n}^{*}\right)-\log S_{n}\left(b_{n}\right)\right)\leq\left(k-1\right)\log\left(n+1\right)}$ for all $n$, as needed to show.
@@
Then, we have the following bound for every representative strategy:
@@
On the other hand, by Proposition 4 and the property of the benchmark strategy for this context, namely $S_{n}\big{(}b_{n}^{*}\big{)}\geq\max_{i\in\left[N\right]}S_{n}\big{(}\hat{b}_{n}^{i}\big{)}$ for all $n$, we obtain the needed upper bound as follows:
@@
The last inequality follows from (A.1) and the fact that $S_{n}\big{(}\ddot{b}_{n}\big{)}=\max_{i\in\left[N\right]}S_{n}\big{(}b_{n}^{\alpha_{n}^{i}}\big{)}$ by definition.